# Korteweg-de Vries Integrals for Modified Black Hole Potentials: Instabilities and other Questions

### v1

@@ 1 Introduction
Recent studies of the spectrum of BH QNMs that use the concept of pseudospectrum indicate that this spectrum could be unstable under small changes in the BH potential [70, 71], which has generated an active area of research (see, e.g. [72, 73, 74, 75, 76]). These studies have carried beyond the QNM spectrum into the greybody factors and, in particular, their stability under modifications of the BH potential has been considered recently [77, 78, 79, 80]. Due to the relevance of these results for the understanding of BH dynamics and the implications for GW physics, in this work we have looked at these questions from a different point of view, using the machinery of hidden symmetries of BH dynamics described above. In particular, we analyze in a number of situations whether the KdV conserved quantities can be used as an indicator of the appearance of instabilities when the BH potential is modified due to different reasons. The different scenarios we have considered are: (i) The Pöschl-Teller potential, for which the Schrödinger equation can be integrated in closed form, fitted to the standard BH potentials. This fit is well-known to provide a good approximation for the BH QNM modes in the eikonal limit. (ii) Astrophysically-motivated environmental corrections to the BH potential. (iii) Potentials with small-scale (oscillatory) modifications. (iv) Potentials coming from Effective Field Theory (EFT) modifications of GR. For all these cases, we analyze in details the properties of the KdV integrals as we modify the different physical parameters that characterize these modified potentials. In particular, the size of the modifications, their main location, and their width. The outcome of this study is that KdV integrals are a very interesting tool to characterize and predict the possibility of introducing instabilities with the modification of the BH potentials. Furthermore, the importance of KdV integrals in BH spectral studies is even clearer once one realizes that the BH greybody factors are completely and uniquely determined by the KdV integrals, as shown in Refs. [56, 57]. This gives a solid mathematical handle to investigate the stability of greybody factors. While our study is of a phenomenological character, it motivates future more in-depth studies.
@@
Structure of the Paper. In Section 2 we introduce the basic elements of BHPT that we need in this work. In particular, the structure of the space of master functions and equations (Sec. 2.1), how isospectrality arises from Darboux covariances (Sec. 2.2), and the emergence of integrable structures that lead to the KdV symmetries and conserved quantities (Sec. 2.3). In Sec. 3 we explain how the KdV integrals completely determine the BH greybody factors via a moment problem. In Sec. 4, we fit the BH potentials (here we consider both the Regge-Wheeler and Zerilli potentials for odd- and even-parity perturbations, respectively) to the Pöschl-Teller potential, as an approximation to the exact potentials (Sec. 4.1). Then, we study some criteria, based on the KdV integrals, to characterize how good as an approximation are these fits (Sec. 4.2). Finally, we compare the previous criteria with the computation of QNMs both for the fits and comparing with the actual values (Sec. 4.3). In Sec. 5 we compute the KdV integrals for modified BH potentials with the aim of studying the connection between stability under these modifications and the behaviour of the KdV integrals. We consider different classes of BH potentials: Potentials that represent (astrophysical) environmental modifications (Sec. 5.1); phenomenological oscillatory potentials (Sec. 5.2); and potentials coming from EFT modifications of GR (Sec. 5.3). In Sec. 6 we discuss the question of stability in terms of the greybody factors. Sec. 7 is devoted to conclusions and discussions of the future prospects of this work. We have some appendices where we include some details of the scattering problem for the Pöschl-Teller potential, for which the master equation in frequency domain is exactly solvable (Appendix A). In Appendix B we provide details for the derivation of the general formula for KdV integrals for the Pöschl-Teller potential shown in Sec. 4.2. In Appendix C we present general and simplied expressions for the KdV densities by using essentially the integration by parts method. Finally, in Appendix D, we include the values of the coefficients that determine the BH potentials from the EFTs of Sec. 5.3. Throughout this paper, otherwise stated, we use geometric units in which $G=c=1\,$.
@@ 2 Perturbed BHs: Master Equations and Hidden Symmetries
where $r_{s}$ is the Schwarzschild radius, $r_{s}=2GM/c^{2}=2M\,$, providing the location of the event horizon. One of the main advantages of BH perturbation theory resides in the possibility of decoupling the linearized Einstein equations for the perturbations into an infinite set of master equations. Exploiting the spherical symmetry of the background, one can decompose every perturbative quantity in scalar, vector and tensor harmonics [^1] These terms may have different meaning depending on the context, e.g. when studying perturbations in higher dimensions [86]. so that the Einstein equations decouple for each harmonic $(\ell,m)$ and parity [^2] Under parity transformation $(\theta,\phi)$rightarrow $(\pi-\theta,\phi+\pi)$ (i.e. spatial inversion), an even-parity harmonic ${\cal O}^{\ell m}$ transforms as ${\cal O}^{\ell m}\rightarrow(-1)^{\ell}{\cal O}^{\ell m}$, while an odd-parity one transforms as ${\cal O}^{\ell m}\rightarrow(-1)^{\ell+1}{\cal O}^{\ell m}$.. Manipulations of such equations can lead to further simplification (thanks also to variable separation), so that the Einstein equations for the perturbations, for each harmonic and parity, reduce to a simple master equations (see, e.g. [42, 87, 58]):
@@ 3 BH Greybody Factors and Korteweg-de Vries Integrals
$\psi(x,k)=\left\{\begin{array}[]{lc}a(k)e^{ikx}+b(k)e^{-ikx}&\mbox{for~{}}x\to-\infty\,,\\[14.22636pt] e^{ikx}&\mbox{for~{}}x\to\infty\,,\end{array}\right.$ (3.3)
@@ 4.1 Fits to the Pöschl-Teller Barrier
Figure 1: Regge-Wheeler (blue) and Zerilli (red) potentials as functions of the tortoise coordinate $x$ ($-\infty<x<\infty$) in units of the Schwarzschild radius $r_{s}$, for angular momentum harmonic number $\ell=2$.
@@ 4.2 KdV Integrals for the Pöschl-Teller Potential: Application to the BH Potential Fits
$+$ $\sum_{m=0}^{n-1}\frac{(2\beta)^{2n-2m}(-1)^{m+n}(2n-2)!(2m-1)}{(2m)!(2n-2m)!}\,,\quad n=1,2,3...$
@@ 4.3 Quasinormal Modes associated with the Pöschl-Teller Fits
Comparing the left and right panels of Figs. 11 and 12, we can notice the improvement of the Pöschl-Teller approximation when $\ell$ grows, as dictated by the eikonal limit. In addition, considering high values of $\ell$ discloses an asymmetry between the two fits which could not have been noticed at lower angular momentum values. In fact, while the two approximations are almost equivalent for the real part of the QNMs (see right panel of Fig. 11), the fit to the Regge-Wheeler potential provides significantly better results for the imaginary parts. For example, the error at $n=60$ is $5$% for the Regge-Wheeler fit while it is around $11$% for the Zerilli fit (see the right panel of Fig. 12). The final qualitative statement that can be done is that the two fits to a Pöschl-Teller potential break the isospectrality of the QNMs.
@@ 5.1 Environmental Corrections to the BH Potentials
[…]
@@
[…]
@@ 5.3 BH Potentials from Effective Field Theory Modifications of GR
Figure 22: Relative error $\delta\mathcal{K}_{2n-1}$ (5.2) for $n=1,...,30$ for the odd (blue dots) and even (black dots) potential and the line corresponding to the value of $\epsilon$ (orange dashed line). In this case $\epsilon=10^{-5}$. The vertical axis is in logarithmic scale of base $10$.
@@
Figure 23: Relative error $\delta\mathcal{K}_{1}$ (left panel) and $\delta\mathcal{K}_{7}$ (5.2) (right panel) for $\epsilon$ ranging from $10^{-15}$ to $2$ for odd (blue dots) and even (black dots) parity. The diagonal line corresponds to plotting $\epsilon$ vs $\epsilon$ (orange line). The functions are in logarithmic scale of base $10$.
@@ 6 Stability of the Greybody Factors: A Moment Problem Perspective
[…]
@@
The recent claims that ringdown signals may be encoded, at least partially, in the greybody factors [77, 78], in addition to the frequently used sum of QNMs oscillations, enrich the discussion on the relation between the QNMs instability and the stability of the GW signal in the time domain. This has then led to consider stability properties of the greybody factors [79, 80] in contrast to those of the QNMs ones [70, 71]. Our results on the instability of the KdV integrals under perturbations opens the door to a systematic study of the stability of the greybody factors within the framework of the moment problem introduced in Refs. [56, 57] and briefly reviewed in Sec. 3. While such formal and analytic study inevitably requires a completely new investigation, which we postpone to future developments, we can already introduce some elements of analysis and prediction, serving as a guide for future works on the greybody factor stability. In fact, a simple intuition can be obtained from the first greybody factor approximants defined in Eq. (3.12). Indeed, let us first recall the analytic expressions for the greybody factors obtained in Ref. [57], in particular those coming from the $[0/1]$ and $[1/1]$ Padé approximants:
@@
[…]
The values of this integrated error are plotted in Fig. 26 for different approximants. As we can see, they are smaller than $\epsilon=10^{-5}$, as required. It is not clear whether such behavior is to be attributed to the breaking of our approximation for very small frequencies (see Ref. [57]) or to the fact that higher KdV integrals tend to be more unstable as we have seen in this work. Moreover, let us stress that a full stability assessment of these quantities requires a deeper numerical and analytical study, which definitely deserves more extensive investigations. Nevertheless, the results we have shown here can be considered as preliminary results and the motivation for a more detailed analysis in the future that has to take into account the following points: i) The conditions imposed by the moment problem on the moments in order for a solution to exist and be unique; ii) higher orders in the approximation to test the convergence properties, since the two approximations considered here are only qualitatively good; iii) analytical tools in the context of the moment problem.

### v3

@@ 1 Introduction
Recent studies of the spectrum of BH QNMs that use the concept of pseudospectrum indicate that this spectrum could be unstable under small changes in the BH potential [70, 71], which has generated an active area of research (see, e.g. [72, 73, 74, 75, 76]). These studies have carried beyond the QNM spectrum into the greybody factors and, in particular, their stability under modifications of the BH potential has been considered recently [77, 78, 79, 80, 81]. Due to the relevance of these results for the understanding of BH dynamics and the implications for GW physics, in this work we have looked at these questions from a different point of view, using the machinery of hidden symmetries of BH dynamics described above. In particular, we analyze in a number of situations whether the KdV conserved quantities can be used as an indicator of the appearance of instabilities when the BH potential is modified due to different reasons. The different scenarios we have considered are: (i) The Pöschl-Teller potential, for which the Schrödinger equation can be integrated in closed form, fitted to the standard BH potentials. This fit is well-known to provide a good approximation for the BH QNM modes in the eikonal limit. (ii) Astrophysically-motivated environmental corrections to the BH potential. (iii) Potentials with small-scale (oscillatory) modifications. (iv) Potentials coming from Effective Field Theory (EFT) modifications of GR. For all these cases, we analyze in details the properties of the KdV integrals as we modify the different physical parameters that characterize these modified potentials. In particular, the size of the modifications, their main location, and their width. The outcome of this study is that KdV integrals are a very interesting tool to characterize and predict the possibility of introducing instabilities with the modification of the BH potentials. Furthermore, the importance of KdV integrals in BH spectral studies is even clearer once one realizes that the BH greybody factors are completely and uniquely determined by the KdV integrals, as shown in Refs. [56, 57]. This gives a solid mathematical handle to investigate the stability of greybody factors. While our study is of a phenomenological character, it motivates future more in-depth studies.
@@
Structure of the Paper. In Section 2 we introduce the basic elements of BHPT that we need in this work. In particular, the structure of the space of master functions and equations (Sec. 2.1), how isospectrality arises from Darboux covariances (Sec. 2.2), and the emergence of integrable structures that lead to the KdV symmetries and conserved quantities (Sec. 2.3). In Sec. 3 we explain how the KdV integrals completely determine the BH greybody factors via a moment problem. In Sec. 4, we fit the BH potentials (here we consider both the Regge-Wheeler and Zerilli potentials for odd- and even-parity perturbations, respectively) to the Pöschl-Teller potential, as an approximation to the exact potentials (Sec. 4.1). Then, we study some criteria, based on the KdV integrals, to characterize how good as an approximation are these fits (Sec. 4.2). Finally, we compare the previous criteria with the computation of QNMs both for the fits and comparing with the actual values (Sec. 4.3). In Sec. 5 we compute the KdV integrals for modified BH potentials with the aim of studying the connection between stability under these modifications and the behaviour of the KdV integrals. We consider different classes of BH potentials: Potentials that represent (astrophysical) environmental modifications (Sec. 5.1); phenomenological oscillatory potentials (Sec. 5.2); and potentials coming from EFT modifications of GR (Sec. 5.3). In Sec. 6 we discuss the question of stability in terms of the greybody factors. Sec. 7 is devoted to conclusions and discussions of the future prospects of this work. We have some appendices where we include some details of the scattering problem for the Pöschl-Teller potential, for which the master equation in frequency domain is exactly solvable (Appendix A). In Appendix B we provide details for the derivation of the general formula for KdV integrals for the Pöschl-Teller potential shown in Sec. 4.2. In Appendix C we present general and simplied expressions for the KdV densities by using essentially the integration by parts method. Finally, in Appendix D, we include the values of the coefficients that determine the BH potentials from the EFTs of Sec. 5.3. Throughout this paper, otherwise stated, we use geometric units in which G=c=1G=c=1\,.
@@ 2 Perturbed BHs: Master Equations and Hidden Symmetries
where $r_{s}$ is the Schwarzschild radius, $r_{s}=2GM/c^{2}=2M\,$, providing the location of the event horizon. One of the main advantages of BH perturbation theory resides in the possibility of decoupling the linearized Einstein equations for the perturbations into an infinite set of master equations. Exploiting the spherical symmetry of the background, one can decompose every perturbative quantity in scalar, vector and tensor harmonics [^1] These terms may have different meaning depending on the context, e.g. when studying perturbations in higher dimensions [87]. so that the Einstein equations decouple for each harmonic $(\ell,m)$ and parity [^2] Under parity transformation $(\theta,\phi)$ $\rightarrow$ $(\pi-\theta,\phi+\pi)$ (i.e. spatial inversion), an even-parity harmonic ${\cal O}^{\ell m}$ transforms as ${\cal O}^{\ell m}\rightarrow(-1)^{\ell}{\cal O}^{\ell m}$, while an odd-parity one transforms as ${\cal O}^{\ell m}\rightarrow(-1)^{\ell+1}{\cal O}^{\ell m}$.. Manipulations of such equations can lead to further simplification (thanks also to variable separation), so that the Einstein equations for the perturbations, for each harmonic and parity, reduce to a simple master equations (see, e.g. [42, 88, 58]):
@@ 3 BH Greybody Factors and Korteweg-de Vries Integrals
[…]
@@ 4.1 Fits to the Pöschl-Teller Barrier
Figure 1: Regge-Wheeler (blue) and Zerilli (red) potentials as functions of the tortoise coordinate $x$ ($\infty<x<\infty$) in units of the Schwarzschild radius $r_{s}$, for angular momentum harmonic number $\ell=2$.
@@ 4.2 KdV Integrals for the Pöschl-Teller Potential: Application to the BH Potential Fits
$+$ $\sum_{m=0}^{n-1}\frac{(2\beta)^{2n-2m}(-1)^{m+n}(2n-2)!(2m-1)}{(2m)!(2n-2m)!}\,,\quad n=1,2,3,\ldots$
@@ 4.3 Quasinormal Modes associated with the Pöschl-Teller Fits
Comparing the left and right panels of Figs. 11 and 12, we can notice the improvement of the Pöschl-Teller approximation when $\ell$ grows, as dictated by the eikonal limit. In addition, considering high values of $\ell$ discloses an asymmetry between the two fits which could not have been noticed at lower angular momentum values. In fact, while the two approximations are almost equivalent for the real part of the QNMs (see right panel of Fig. 11), the fit to the Regge-Wheeler potential provides significantly better results for the imaginary parts. For example, the error at $n=60$ is 55% for the Regge-Wheeler fit while it is around 1111% for the Zerilli fit (see the right panel of Fig. 12). The final qualitative statement that can be done is that the two fits to a Pöschl-Teller potential break the isospectrality of the QNMs.
@@ 5.1 Environmental Corrections to the BH Potentials
$\mathcal{K}^{\epsilon}_{3}=\mathcal{K}^{\mathrm{RW}}_{3}+\mathcal{K}_{3}[\epsilon\delta V]-2\epsilon\int^{+\infty}_{-\infty}dx\,V^{\mathrm{RW}}\delta V\,.$ (5.10)
@@
$\int^{+\infty}_{-\infty}dx\,V^{\mathrm{RW}}\delta V\approx V^{\mathrm{RW}}(x_{0})\int^{+\infty}_{-\infty}dx\,\delta V\approx\frac{\ell(\ell+1)}{r^{2}_{0}}\frac{2}{r_{s}\hat{\alpha}}\,$. (5.11)
@@ 5.3 BH Potentials from Effective Field Theory Modifications of GR
Figure 22: Relative error $\delta\mathcal{K}_{2n-1}$ (5.2) for n=1,,30n=1,...,30 for the odd (blue dots) and even (black dots) potential and the line corresponding to the value of $\epsilon$ (orange dashed line). In this case $\epsilon=10^{-5}$. The vertical axis is in logarithmic scale of base $10$.
@@
Figure 23: Relative error $\delta\mathcal{K}_{1}$ (left panel) and $\delta\mathcal{K}_{7}$ (5.2) (right panel) for $\epsilon$ ranging from $10^{-15}$ to 22 for odd (blue dots) and even (black dots) parity. The diagonal line corresponds to plotting $\epsilon$ vs $\epsilon$ (orange line). The functions are in logarithmic scale of base $10$.
@@ 6 Stability of the Greybody Factors: A Moment Problem Perspective
$\mathcal{K}_{2n+1}=\int^{\infty}_{-\infty}dx\,\kappa_{2n+1}(x)=(-1)^{n+1}2^{2n+1}\int^{\infty}_{-\infty}dk\,k^{2n}\frac{\ln{T(k)}}{2\pi}\,$, (6.1)
@@
The moment problem perspective, see Eqs. (3.8) and (3.9), already gives a first insight on the stability of the logarithmic distribution of the greybody factors. Indeed, as shown in Fig. 24, the $\ln{T}$ distribution is mostly stable (within one order of magnitude) for frequencies below (the real part of) the Schwarzschild fundamental QNM frequency  [^8] The oscillations in the low frequency limit are accounted by the approximation procedure [see Eq. (3.12)] and improve when increasing the order of the Padé approximants [57].. The instability becomes relevant for higher frequencies. This behavior perfectly reflects the fact that the lower order KdV integrals/moments tends to be (almost) stable even for narrow bump perturbations, while the higher KdV integrals/moments becomes strongly unstable in this case (see Fig. 15), thus explaining the instability/stability of the distribution at high/low frequencies.
Figure 24: Plot of the relative error on the distribution $\ln{T_{[N/M]}}$ (left panel) and the corresponding logarithmic plot (right panel) evaluated with the approximation in Eq. (3.12) for the Pöschl-Teller bump correction (5.4) for $\alpha=1$ (narrow bump), $x_{0}=80r_{s}$ and $\epsilon=10^{-2}$. The black dashed lines correspond to the limiting values $\epsilon=10^{-2}$ and $\epsilon=10^{-1}$. The black dotted line marks the value of the real part of the Schwarzschild QNM frequency. Note that we only plot the subdiagonal Padé approximants ($M=N+1$) because the qualitative behavior is the same for the diagonal ones ($M=N$).
The recent claims that ringdown signals may be encoded, at least partially, in the greybody factors [77, 78, 79], in addition to the frequently used sum of QNMs oscillations, enrich the discussion on the relation between the QNMs instability and the stability of the GW signal in the time domain. This has then led to consider stability properties of the greybody factors [80, 81] in contrast to those of the QNMs [70, 71]. Our results on the instability of the KdV integrals under perturbations opens the door to a systematic study of the stability of the greybody factors within the framework of the moment problem introduced in Refs. [56, 57] and briefly reviewed in Sec. 3. While such formal and analytic study inevitably requires a completely new investigation, which we postpone to future developments, we can already introduce some elements of analysis and prediction, serving as a guide for future works on the greybody factor stability. In fact, a simple intuition can be obtained from the first greybody factor approximants defined in Eq. (3.12). Indeed, let us first recall the analytic expressions for the greybody factors obtained in Ref. [57], in particular those coming from the $[0/1]$ and $[1/1]$ Padé approximants:
@@
$\Delta T_{[N/M]}=\frac{\int_{0}^{\infty}dk\left|T^{\mathrm{bump}}_{[N/M]}-T^{\mathrm{RW}}_{[N/M]}\right|}{\int_{0}^{\infty}dk\,T^{\mathrm{RW}}_{[N/M]}}\,.$ (6.7)
The values of this integrated error are plotted in Fig. 27 for different approximants. As we can see, they are smaller than $\epsilon=10^{-5}$, as required. It is not clear whether such behavior is to be attributed to the breaking of our approximation for very small frequencies (see Ref. [57]) or to the fact that higher KdV integrals tend to be more unstable as we have seen in this work.
In relation to this, in Refs. [77, 78, 79] the ringdown spectral amplitude is proportional to the reflection coefficient (reflectivity), i.e.
$|h_{\ell,m}(k)|\propto\sqrt{1-T_{\ell,m}(k)}=\sqrt{R_{\ell,m}(k)}\,.$ (6.8)
It is therefore instructive to consider the reflectivity $R_{[N/M]}=1-T_{[N/M]}$, facilitating a better comparison with the results of Refs. [80, 81]. In Fig. 28 we show the relative error of the $R_{[N/M]}$. This plot reveals an instability of the reflectivity for frequencies above the real part of the Schwarzschild fundamental QNM, in accordance with the results of Ref. [80]. This is known to be related to the fact that the reflectivity is exponentially suppressed at high frequencies [80, 81] so the relative error ends up growing in this regime. In particular, the exponential decay is governed by the WKB approximation for the Schwarzschild case, i.e.
$R\sim e^{-\kappa_{0}k}\,,\quad\kappa_{0}=2\pi\sqrt{\frac{2V_{\rm RW}(x_{*})}{-V_{\rm RW,xx}(x_{*})}}\,$, (6.9)
where $x_{*}$ is the maximum of the Regge-Wheeler potential in the tortoise coordinate. In Ref. [80] the authors further showed that high frequency limit of the reflectivity for perturbed potentials follows a similar exponential decay but with the coefficient in the exponential fixed by the maximum at the bump position, $x=x_{0}$ in our notation (see Sec. 5.1). In other words:
$R\sim e^{-\kappa_{\delta}k}\,,\quad\kappa_{\delta}=2\pi\sqrt{\frac{2V(x_{0})}{-V_{,xx}(x_{0})}}\,$, (6.10)
where $V$ is the potential given in Eq. (5.1) with $\delta V$ given by Eq. (5.4) and $x_{0}$ is the maximum of the bump $\delta V$. Fig. 29 (left panel) actually confirms this behavior as we observe that the approximation correctly reproduce the exponential decay (6.10) at high frequency. Of course, the exact asymptotic behavior can only be obtained in the limit of infinite KdV integrals (the convergence of the approximation is studied in Ref. [57]). The right panel of Fig. 29 shows instead the plot for $\ln{T}$, which is proportional to the moment problem distribution (see Eq. (3.9)). It is quite interesting to note that the high frequency behavior seems to match the one of the reflectivity with the same asymptotic decay. This is because, for high frequency we have
$\ln{T}=\ln{(1-R)}\simeq\ln{\left(1-e^{-\kappa k}\right)}\simeq e^{-\kappa k}\simeq R\,.$ (6.11)
Taking into account that the instability behavior of $\ln{T}$ is understood in terms of the instability of the higher KdV integrals for narrow bumps (see Fig. 24), this seems to suggest a strong relation between the moment problem perspective and the instability of the reflectivity R=1TR=1-T  [^9] We thank an anonymous referee for pointing out Eq. (6.11) in relation to this.. However, let us stress that a full stability assessment of these quantities requires a deeper numerical and analytical study, which definitely deserves more extensive investigations. Nevertheless, the results we have shown here can be considered as preliminary results and the motivation for a more detailed analysis in the future that has to take into account the following points: i) The conditions imposed by the moment problem on the moments in order for a solution to exist and be unique; ii) higher orders in the approximation to test the convergence properties, since the two approximations considered here are only qualitatively good; iii) analytical tools in the context of the moment problem.
@@
Figure 28: Plot of the relative error on $R_{[N/M]}$ (left panel) and the corresponding logarithmic plot (right panel) evaluated with the approximation in Eq. (3.12) for the Pöschl-Teller bump correction (5.4) for $\alpha=1$ (narrow bump), $x_{0}=80r_{s}$ and $\epsilon=10^{-2}$. The black dashed lines in the left panel correspond to the limiting values $\epsilon=10^{-2}$ and $\epsilon=10^{-1}$. The black dotted line marks the value of the real part of the Schwarzschild QNM frequency. Again, we only plot the subdiagonal Padé approximants ($M=N+1$) because the qualitative behavior is the same for the diagonal ones ($M=N$).
Figure 29: Plot of $R_{[N/M]}$ (left panel) and $\ln{T_{[N/M]}}$ (right panel) evaluated with the approximation in Eq. (3.12) for the Pöschl-Teller bump correction (5.4) for $\alpha=1$ (narrow bump), $x_{0}=80r_{s}$ and $\epsilon=10^{-2}$. Here $c_{\delta}=0.0014\,,r_{s}\kappa_{\delta}=6.59$ and $c_{0}=10^{8}\,,r_{s}\kappa_{0}=17.35$. We only plot the subdiagonal Padé approximants ($M=N+1$) because the qualitative behavior is the same for the diagonal ones ($M=N$).